Trying to help my child and I do not understand how to solve these problems.
Equations as follows:
y= x-1
y= -x +7


Can you please show solution with explanation so I understand where the answers (numbers) came from and how these problems are solved?

Answer by scott8148(6628)

with two equations and two unknowns, there are various methods of solution

in this case, the easiest is to add the equations

y= x-1
y= -x +7
_________
2y = 6

so, y=3 ____ substituting ___ 3 = x-1 ___ so, x=4




another technique is to substitute at the start ___ in this case, for the y value

x - 1 = -x + 7 ___ 2x = 8 ___ again, x = 4

then substitute to find y ___ y = 4 - 1 ___ y = 3



the choice of the method of solution is usually a combination of the complexity of the problem and personal preference